# ARCANE: Adaptive Routing with Caching and Aware Network Exploration

### v4

@@ 4. Evaluation
Our evaluation includes both large-scale simulations designed to stress test ARCANE under a variety of workloads, and experiments on real hardware at meaningful scale. The goal is to answer the following research questions:
  • Does ARCANE outperform state-of-the-art load balancers in a healthy network with a symmetric topology? (Sections 4.3.1 and 4.4.1)
  • How does ARCANE perform in the presence of topology asymmetries or when coexisting with non-ARCANE traffic? (Sections 4.3.2 and 4.4.2)
  • Can ARCANE recover quickly from failures? Is it resilient to severe or widespread failures? (Sections 4.3.3 and 4.4.3)
  • Does ARCANE maintain performance under varying network parameters and configurations? (Section 4.5)
@@ 4.1. Evaluation Setup
In large-scale simulations, we compare ARCANE with ECMP (Hopps, 2009), OPS (Dixit et al., 2013), PLB (Qureshi et al., 2022), MPRDMA (Lu et al., 2018), Flowlet Switching (Vanini et al., 2017), MPTCP (Raiciu et al., 2011), a bitmap approach where we keep per EV statistics similarly to STrack (Le et al., 2024), and adaptive RoCE by NVIDIA (NVIDIA, 2024). We note that this list includes both sender-based approaches and in-network switch-based approaches. We configure PLB to have more aggressive parameters similarly to FlowBender to improve its performance (Kabbani et al., 2014). For Flowlet Switching we set an aggressive flowlet timeout at half of the RTT. For MPTCP, we divide each message into 8 subflows and route each one individually similarly to using multiple QPs (Queue Pairs) (Gangidi et al., 2024).
@@
Regardless of the workloads that we discuss in Section 4.2, in the evaluation we simulate 3 different scenarios: (1) healthy symmetric network conditions, (2) asymmetric network conditions, which can be due to asymmetric path bandwidth or due to ECMP-hashed background traffic that increase load on specific paths (a scenario that can happen in incremental deployments of ARCANE), and (3) a network encountering various failures modes. We focus on the most relevant failure scenarios for real-world deployments (Gangidi et al., 2024) and we report the remaining ones in Appendix D.
@@
We implement ARCANE by extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider different fat-tree topologies with 1024 nodes and 128 nodes and with different levels of oversubscription ranging from 1:1 (no oversubscription) to 4:1. We test 2- and 3-tier fat trees (TOR or Top-of-rack as T0, Aggregate as T1 and Core T2). Such topologies are commonly deployed in production datacenters designed for distributed training (Singh et al., 2015; Gangidi et al., 2024).
@@
ARCANE-FPGA::
To demonstrate the effectiveness and resilience of entropy recycling in a real network environment, we also evaluate ARCANE in an end-to-end setting. We use a modified production-grade FPGA-based RDMA-capable NIC running a custom transport protocol capable of handling OOO packets by directly placing the payload in memory via RDMA. The transport implementation relies on SACKs that carry bitmaps (256-bit-wide in these experiments) relaying packet delivery updates to track completions and handle selective retransmissions. Each connection maintains an 8-entry-deep ARCANE buffer that stores UDP source ports (2B). Our experimental setup is configured to support up to 256 connections for a total memory footprint of 4KB. Given that only a single connection is active at a time, all ARCANE buffers can be hosted in a single SRAM memory and the associated logic to manage ARCANE buffers can be multiplexed across all connections, occupying negligible resources (¡0.04% of the total device logic resources).
@@ 4.2. Workloads
We evaluate ARCANE on a mix of synthetic benchmarks, real datacenter traces, and distributed training collectives.
@@
Our baseline traffic for ARCANE-FPGA consists of 128 endpoints under two T0 switches continuously performing 4 MB ring-based AllReduce collective operations, with the logical ring laid out such that all connections traverse the T1 spine to maximize the pressure on the spine of the topology.
@@ 4.3. Simulation Results
We conduct a detailed analysis of ARCANE behavior for each network condition (see Section 4). We first examine a specific case in depth and then summarize key takeaways.
@@ 4.3.1. Healthy Symmetric Network Conditions
In this Section we evaluate the performance of ARCANE in a simple setting where there is no oversubscription and there are no failures, meaning the network is perfectly symmetrical. Intuitively, this seems the best situation for oblivious packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we will see later in this section and based on our theoretical model in Section 5, this is not the case as ARCANE still offers an up to 25% advantage over OPS. This is because of ECMP collisions that still happen with OPS. While over long period of time, each link will be evenly used, there will still be short-term collisions happening that will increase and decrease the link utilization of certain links.
Figure 1. A tornado workload with 16 MiB messages using OPS (top) and ARCANE (bottom) as load balancers.
Figure 2. ARCANE performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives.
Figure 3. ARCANE vs. OPS in a 32 MiB message send with an asymmetric topology.
@@
In the same configuration with ARCANE, we notice a major difference in the bottom plot of Figure 1. In particular ARCANE converge quickly to a configuration where each queue is kept below $K_{min}$ (note that the only guarantee here is that all queues will be below $K_{min}$, not necessarily all at the same value). At the same time, we can also notice that all the ports converge to the perfect selection rate of 400 Gbps. While the overall completion time is only about 4% better than OPS, the smaller queues provide a better guarantees for system low latency traffic. Moreover, as we can see in Figure 2, this gap expands as we increase the message size.
Looking at the port selection rate over the entire run of the workload for OPS vs. ARCANE, we observe that they are nearly equivalent. This is, again, because the problem is with the short-term collisions that OPS can experience at microscopic scale.
@@
We now focus more on the overall view comparing ARCANE with all the other state of the art algorithms in a series of benchmarks.
In Figure 2, we visualize a summary of the performance of the various algorithms by looking at the runtime of the workloads (max FCT). As expected, in the case of incast, the performance is driven almost exclusively by the CC and, hence, we do not see any major difference between all the load balancers and even ECMP performs well. However, once we move to permutation and tornado workloads, ECMP collisions start to drastically reduce the performance of ECMP. In most cases ARCANE outperforms all the other algorithms.
In the tornado case, Adaptive RoCE is able to match ARCANE since this is the ideal scenario for it: ARCANE, unlike Adaptive RoCE, still needs to guess during its initial BDP worth of packets. On the other hand, ARCANE outperforms Adaptive RoCE in the permutation pattern where taking a local optimal decision might not always lead to the best global outcome. We also see the difference between algorithms that were designed to reduce the number of out-of-order packets versus algorithms that do not have such hard constraints. Additionally, there is a distinction between algorithms that operate at packet-level granularity, such as ARCANE, OPS, BitMap, and MPRDMA, and those that operate at a coarser granularity, such as Flowlet and PLB.
For datacenter traces, we analyze the results for different load levels, ranging from 40% to 100%. Here we observe again, the clear difference between per-packet algorithms and less granular options. Even at higher load ARCANE is able to work well with a 5% advantage over OPS.
We also run distributed collectives and report their completion times. We observe how, by design, the ring AllReduce has the same performance for most of the load balancing algorithms utilized. This is because, due to its ring design, there is no opportunity for congestion to accumulate. In AllToAll, ARCANE gets an up to 20% advantage over the alternatives.
Figure 4. ARCANE performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives and an asymmetric network due to 2% of the TOR uplinks being offline.
Figure 5. Synthetic traffic with ARCANE and non-ARCANE (ECMP) traffic, emulating a mixed traffic scenario.
@@ 4.3.2. Asymmetric Network Conditions
We evaluate ARCANE under different scenarios where some degree of asymmetry is created. We focus on two scenarios: 1) the network is not perfectly symmetrical because of some missing (or degraded) cables, 2) there is some background traffic in the network that is using ECMP routing, analyzing the coexistence between ARCANE and non-ARCANE traffic.
@@
We visualize this problem with a simple scenario where we have a switch with $n$ input and output ports and $n$ flows active, each from a different source sending a 32 MiB message. To create an asymmetry, we reduce one of the uplinks speed to 200 Gbps while all the other links remain at 400 Gbps. In Figure 3 we visualize the output port utilization rate for OPS and ARCANE.
We observe that while OPS chooses each port equally, irrespective of its actual bandwidth, ARCANE eventually converges to a stable configuration where the slower uplink is used less frequently. This results in both stabler queues but, more importantly, a much faster completion time ($1400\;\mu\text{s}$ for OPS and $756\;\mu\text{s}$ for ARCANE).
@@
We now shift our focus to more general results when encountering asymmetries in a network. For space constraints, we focus mostly on the case where some of the links have a lower sending rate. In our first experiment we run synthetic benchmarks where $2\%$ of the TOR uplinks, chosen randomly, have been downgraded to 200 Gbps. In Figure 4, we can see results similar to before where ARCANE gets an up to 350% advantage over ECMP and 10% advantage over the second best algorithm (usually BitMap). In the DC traces we can see a higher difference due to the asymmetry in the network. At 100% load ARCANE gets a 25% advantage over the second best algorithm and a 1000% advantage over ECMP. We show the results for several AI collectives. We note how for AllToAll ARCANE keep a small but significant advantage and in the AllReduce a sizable 50% advantage over the second best performing algorithm.
In Figure 5, we showcase one example of ARCANE sharing traffic together with background ECMP traffic (we assume 10% of the traffic is ECMP). In this case, ARCANE: 1) shifts ARCANE traffic away from ECMP traffic in order to not slow down ARCANE traffic, 2) helps background traffic by ensuring that it will not be slowed down by ARCANE traffic. This also highlights the possibility of incrementally deploying ARCANE on ECMP-base systems. Finally, we note that WCMP (Zhou et al., 2014) could be used to enhance ECMP performance in the case of a topology with known asymmetries, but would not help as much in the case of unpredictable mixed traffic or sudden temporary asymmetries.
Figure 6. ARCANE vs. OPS in a 32 MiB permutation with two cables’ failure (a shorter one and a longer one).
Figure 7. ARCANE performance under different failure modes in a 8 MiB permutation, DC traces at 100% load and a ring AllReduce.
@@ 4.3.3. Network Failures
In Figure 6, OPS keeps choosing all paths equally (although at lower rate due to CC activation), while ARCANE, once it enters freezing mode, stops selecting the failing paths all together after only one timeout period (order of tens of microseconds). Afterwards, once the failure stops, ARCANE also exits freezing mode and converges once again quickly to use all paths. The overall result is that, compared to OPS, ARCANE completes the workload more than 35% faster even with such a short failure and, more importantly, reduces the number of dropped packets by 2.5×2.5$\times$2.5 ×. We showcase a similar analysis but for incremental and concurrent failures in Appendix D.3.
@@
We showcase three cases with a series of failure modes in Figure 7. We introduce each failure mode after a fixed interval to highlight how ARCANE dynamically reacts to changing network conditions, and then let the failure persist for the remainder of the experiment. We can see when dealing with total failures that ARCANE provides a dramatic speedup over OPS but also other load balancer algorithms. This is because of freezing mode which helps ARCANE stick to the safe path configuration in its cache, allowing it to react immediately upon failure, much faster than the time ECMP routing takes to converge and exclude the failing path. Positively, we note that the gains with ARCANE increase with the amount of failures. Furthermore, random drops do not affect negatively ARCANE performance.
@@
To further demonstrate the resilience of ARCANE, we evaluate its performance under an extreme and unlikely scenario characterized by increasingly large and long-lasting network failures during a permutation. As shown in Figure 8, ARCANE performs close to an ideal load balancer, even with 50% of network cables failing, while PLB, the second-best alternative, lags significantly behind.
This analysis proves that ARCANE is able to guarantee good load balancing performance under short-lived failures, long-lived failures and a combination of the two even with catastrophically high failure rates.
4.4. ARCANE-FPGA Evaluation
@@ 4.4.1. Healthy Symmetric Network Conditions
The first set of results focuses on a baseline healthy symmetric network configuration. Figure 9(a) shows per-flow goodput defined as end-to-end useful bit rate observed by application, after header, overheads, retransmissions, etc. We use OPS and ARCANE across two experimental configurations: denoted as setup-1 and setup-2. In setup-1, all FPGA endpoints under the two T0s are active while in setup-2, 40 out of 64 FPGA endpoints are active.
We present results from both of these configurations as we observe small unexpected performance variations depending on which and how many switch ports are active in experiments near the peak network performance levels using all switch ports (setup-1). These variations appear to be related to internal switch microarchitectural details such as port-buffer affinity and vendor-specific scheduling policies. The setup-2 uses a subset of the switch ports and eliminates most of these vendor-specific and implementation-related behaviors. To get a better understanding of this behavior we performed a sweep where we capped the TX rate of our FPGA NICs and discovered that the slight degradation for setup-1 when using ARCANE appears to subside if the TX rate is capped at 95 Gbps.

### v5

@@ 4. Evaluation
Our evaluation includes both large-scale simulations designed to stress test REPS under a variety of workloads, and experiments on real hardware at meaningful scale. The goal is to answer the following research questions:
  • Does REPS outperform state-of-the-art load balancers in a healthy network with a symmetric topology? (Sections 4.3.1 and 4.4.1)
  • How does REPS perform in the presence of topology asymmetries or when coexisting with non-REPS traffic? (Sections 4.3.2 and 4.4.2)
  • Can REPS recover quickly from failures? Is it resilient to severe or widespread failures? (Sections 4.3.3 and 4.4.3)
  • Does REPS maintain performance under varying network parameters and configurations? (Section 4.5)
@@ 4.1. Evaluation Setup
In large-scale simulations, we compare REPS with ECMP (Hopps, 2009), OPS (Dixit et al., 2013), PLB (Qureshi et al., 2022), MPRDMA (Lu et al., 2018), Flowlet Switching (Vanini et al., 2017), MPTCP (Raiciu et al., 2011), a bitmap approach where we keep per EV statistics similarly to STrack (Le et al., 2024), and adaptive RoCE by NVIDIA (NVIDIA, 2024). We note that this list includes both sender-based approaches and in-network switch-based approaches. We configure PLB to have more aggressive parameters similarly to FlowBender to improve its performance (Kabbani et al., 2014). For Flowlet Switching we set an aggressive flowlet timeout at half of the RTT. For MPTCP, we divide each message into 8 subflows and route each one individually similarly to using multiple QPs (Queue Pairs) (Gangidi et al., 2024).
@@
Regardless of the workloads that we discuss in Section 4.2, in the evaluation we simulate 3 different scenarios: (1) healthy symmetric network conditions, (2) asymmetric network conditions, which can be due to asymmetric path bandwidth or due to ECMP-hashed background traffic that increase load on specific paths (a scenario that can happen in incremental deployments of REPS), and (3) a network encountering various failures modes. We focus on the most relevant failure scenarios for real-world deployments (Gangidi et al., 2024) and we report the remaining ones in Appendix D.
@@
We implement REPS by extending the htsim packet-level network simulator (Handley et al., 2017). Our simulations consider different fat-tree topologies with 1024 nodes and 128 nodes and with different levels of oversubscription ranging from 1:1 (no oversubscription) to 4:1. We test 2- and 3-tier fat trees (TOR or Top-of-rack as T0, Aggregate as T1 and Core T2). Such topologies are commonly deployed in production datacenters designed for distributed training (Singh et al., 2015; Gangidi et al., 2024).
@@
REPS-FPGA::
To demonstrate the effectiveness and resilience of entropy recycling in a real network environment, we also evaluate REPS in an end-to-end setting. We use a modified production-grade FPGA-based RDMA-capable NIC running a custom transport protocol capable of handling OOO packets by directly placing the payload in memory via RDMA. The transport implementation relies on SACKs that carry bitmaps (256-bit-wide in these experiments) relaying packet delivery updates to track completions and handle selective retransmissions. Each connection maintains an 8-entry-deep REPS buffer that stores UDP source ports (2B). Our experimental setup is configured to support up to 256 connections for a total memory footprint of 4KB. Given that only a single connection is active at a time, all REPS buffers can be hosted in a single SRAM memory and the associated logic to manage REPS buffers can be multiplexed across all connections, occupying negligible resources (¡0.04% of the total device logic resources).
@@ 4.2. Workloads
We evaluate REPS on a mix of synthetic benchmarks, real datacenter traces, and distributed training collectives.
@@
Our baseline traffic for REPS-FPGA consists of 128 endpoints under two T0 switches continuously performing 4 MB ring-based AllReduce collective operations, with the logical ring laid out such that all connections traverse the T1 spine to maximize the pressure on the spine of the topology.
@@ 4.3. Simulation Results
We conduct a detailed analysis of REPS behavior for each network condition (see Section 4). We first examine a specific case in depth and then summarize key takeaways.
@@ 4.3.1. Healthy Symmetric Network Conditions
In this Section we evaluate the performance of REPS in a simple setting where there is no oversubscription and there are no failures, meaning the network is perfectly symmetrical. Intuitively, this seems the best situation for oblivious packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we will see later in this section and based on our theoretical model in Section 5, this is not the case as REPS still offers an up to 25% advantage over OPS. This is because of ECMP collisions that still happen with OPS. While over long period of time, each link will be evenly used, there will still be short-term collisions happening that will increase and decrease the link utilization of certain links.
Figure 1. A tornado workload with 16 MiB messages using OPS (top) and REPS (bottom) as load balancers.
Figure 2. REPS performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives.
Figure 3. REPS vs. OPS in a 32 MiB message send with an asymmetric topology.
@@
In the same configuration with REPS, we notice a major difference in the bottom plot of Figure 1. In particular REPS converge quickly to a configuration where each queue is kept below $K_{min}$ (note that the only guarantee here is that all queues will be below $K_{min}$, not necessarily all at the same value). At the same time, we can also notice that all the ports converge to the perfect selection rate of 400 Gbps. While the overall completion time is only about 4% better than OPS, the smaller queues provide a better guarantees for system low latency traffic. Moreover, as we can see in Figure 2, this gap expands as we increase the message size.
Looking at the port selection rate over the entire run of the workload for OPS vs. REPS, we observe that they are nearly equivalent. This is, again, because the problem is with the short-term collisions that OPS can experience at microscopic scale.
@@
We now focus more on the overall view comparing REPS with all the other state of the art algorithms in a series of benchmarks.
In Figure 2, we visualize a summary of the performance of the various algorithms by looking at the runtime of the workloads (max FCT). As expected, in the case of incast, the performance is driven almost exclusively by the CC and, hence, we do not see any major difference between all the load balancers and even ECMP performs well. However, once we move to permutation and tornado workloads, ECMP collisions start to drastically reduce the performance of ECMP. In most cases REPS outperforms all the other algorithms.
In the tornado case, Adaptive RoCE is able to match REPS since this is the ideal scenario for it: REPS, unlike Adaptive RoCE, still needs to guess during its initial BDP worth of packets. On the other hand, REPS outperforms Adaptive RoCE in the permutation pattern where taking a local optimal decision might not always lead to the best global outcome. We also see the difference between algorithms that were designed to reduce the number of out-of-order packets versus algorithms that do not have such hard constraints. Additionally, there is a distinction between algorithms that operate at packet-level granularity, such as REPS, OPS, BitMap, and MPRDMA, and those that operate at a coarser granularity, such as Flowlet and PLB.
For datacenter traces, we analyze the results for different load levels, ranging from 40% to 100%. Here we observe again, the clear difference between per-packet algorithms and less granular options. Even at higher load REPS is able to work well with a 5% advantage over OPS.
We also run distributed collectives and report their completion times. We observe how, by design, the ring AllReduce has the same performance for most of the load balancing algorithms utilized. This is because, due to its ring design, there is no opportunity for congestion to accumulate. In AllToAll, REPS gets an up to 20% advantage over the alternatives.
Figure 4. REPS performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives and an asymmetric network due to 2% of the TOR uplinks being offline.
Figure 5. Synthetic traffic with REPS and non-REPS (ECMP) traffic, emulating a mixed traffic scenario.
@@ 4.3.2. Asymmetric Network Conditions
We evaluate REPS under different scenarios where some degree of asymmetry is created. We focus on two scenarios: 1) the network is not perfectly symmetrical because of some missing (or degraded) cables, 2) there is some background traffic in the network that is using ECMP routing, analyzing the coexistence between REPS and non-REPS traffic.
@@
We visualize this problem with a simple scenario where we have a switch with $n$ input and output ports and $n$ flows active, each from a different source sending a 32 MiB message. To create an asymmetry, we reduce one of the uplinks speed to 200 Gbps while all the other links remain at 400 Gbps. In Figure 3 we visualize the output port utilization rate for OPS and REPS.
We observe that while OPS chooses each port equally, irrespective of its actual bandwidth, REPS eventually converges to a stable configuration where the slower uplink is used less frequently. This results in both stabler queues but, more importantly, a much faster completion time ($1400\;\mu\text{s}$ for OPS and $756\;\mu\text{s}$ for REPS).
@@
We now shift our focus to more general results when encountering asymmetries in a network. For space constraints, we focus mostly on the case where some of the links have a lower sending rate. In our first experiment we run synthetic benchmarks where $2\%$ of the TOR uplinks, chosen randomly, have been downgraded to 200 Gbps. In Figure 4, we can see results similar to before where REPS gets an up to 350% advantage over ECMP and 10% advantage over the second best algorithm (usually BitMap). In the DC traces we can see a higher difference due to the asymmetry in the network. At 100% load REPS gets a 25% advantage over the second best algorithm and a 1000% advantage over ECMP. We show the results for several AI collectives. We note how for AllToAll REPS keep a small but significant advantage and in the AllReduce a sizable 50% advantage over the second best performing algorithm.
In Figure 5, we showcase one example of REPS sharing traffic together with background ECMP traffic (we assume 10% of the traffic is ECMP). In this case, REPS: 1) shifts REPS traffic away from ECMP traffic in order to not slow down REPS traffic, 2) helps background traffic by ensuring that it will not be slowed down by REPS traffic. This also highlights the possibility of incrementally deploying REPS on ECMP-base systems. Finally, we note that WCMP (Zhou et al., 2014) could be used to enhance ECMP performance in the case of a topology with known asymmetries, but would not help as much in the case of unpredictable mixed traffic or sudden temporary asymmetries.
Figure 6. REPS vs. OPS in a 32 MiB permutation with two cables’ failure (a shorter one and a longer one).
Figure 7. REPS performance under different failure modes in a 8 MiB permutation, DC traces at 100% load and a ring AllReduce.
@@ 4.3.3. Network Failures
In Figure 6, OPS keeps choosing all paths equally (although at lower rate due to CC activation), while REPS, once it enters freezing mode, stops selecting the failing paths all together after only one timeout period (order of tens of microseconds). Afterwards, once the failure stops, REPS also exits freezing mode and converges once again quickly to use all paths. The overall result is that, compared to OPS, REPS completes the workload more than 35% faster even with such a short failure and, more importantly, reduces the number of dropped packets by $2.5\times$. We showcase a similar analysis but for incremental and concurrent failures in Appendix D.3.
@@
We showcase three cases with a series of failure modes in Figure 7. We introduce each failure mode after a fixed interval to highlight how REPS dynamically reacts to changing network conditions, and then let the failure persist for the remainder of the experiment. We can see when dealing with total failures that REPS provides a dramatic speedup over OPS but also other load balancer algorithms. This is because of freezing mode which helps REPS stick to the safe path configuration in its cache, allowing it to react immediately upon failure, much faster than the time ECMP routing takes to converge and exclude the failing path. Positively, we note that the gains with REPS increase with the amount of failures. Furthermore, random drops do not affect negatively REPS performance.
@@
To further demonstrate the resilience of REPS, we evaluate its performance under an extreme and unlikely scenario characterized by increasingly large and long-lasting network failures during a permutation. As shown in Figure 8, REPS performs close to an ideal load balancer, even with 50% of network cables failing, while PLB, the second-best alternative, lags significantly behind.
This analysis proves that REPS is able to guarantee good load balancing performance under short-lived failures, long-lived failures and a combination of the two even with catastrophically high failure rates.
4.4. REPS-FPGA Evaluation
@@ 4.4.1. Healthy Symmetric Network Conditions
The first set of results focuses on a baseline healthy symmetric network configuration. Figure 9(a) shows per-flow goodput defined as end-to-end useful bit rate observed by application, after header, overheads, retransmissions, etc. We use OPS and REPS across two experimental configurations: denoted as setup-1 and setup-2. In setup-1, all FPGA endpoints under the two T0s are active while in setup-2, 40 out of 64 FPGA endpoints are active.
We present results from both of these configurations as we observe small unexpected performance variations depending on which and how many switch ports are active in experiments near the peak network performance levels using all switch ports (setup-1). These variations appear to be related to internal switch microarchitectural details such as port-buffer affinity and vendor-specific scheduling policies. The setup-2 uses a subset of the switch ports and eliminates most of these vendor-specific and implementation-related behaviors. To get a better understanding of this behavior we performed a sweep where we capped the TX rate of our FPGA NICs and discovered that the slight degradation for setup-1 when using REPS appears to subside if the TX rate is capped at 95 Gbps.